OpenStudy (anonymous):

rewrite in logarithic form 4^2=16

OpenStudy (anonymous):

\[\log(4^{2})=\log(16)\] \[2*\log(4) = \log(16)\] \[2=\log(16)/\log(4)\] \[2=\log _{4}^{16}\]